OpenStudy (crashonce):

Prove cos 2A + cos 2B + cos 2C + cos 2D = 4[cosAcosBcosCcosD - sinAsinBsinCsinD] If A+B+C+D=0 @ikram002p @ganeshie8 @Hero

OpenStudy (crashonce):

@mayankdevnani

OpenStudy (anonymous):

Cos [ A + B +C +D] = Cos (0) = 1

OpenStudy (anonymous):

Cos [(A + B) + (C+D)] = Cos (A+B)Cos(C+D) - Sin(A+B)Sin(C+D) [Cos A Cos B - Sin A Sin B][Cos C Cos D - Sin C Sin D] - [Sin A Cos B + Cos A Sin B][Sin C Cos D + Cos C Sin D] It might help. just needs to be simplified anyway.

OpenStudy (crashonce):

not sure how that helps @ikram002p @ganeshie8 any ideas?

OpenStudy (midhun.madhu1987):

@lekshmyprasad ? any idea?